OpenStudy (geekfromthefutur):

If the volume for 3-D polyhedra A is 300 cm3 and the volume for 3-D polyhedral B is 900 cm3. How many times bigger is the volume of pyramid B than pyramid A?

OpenStudy (anonymous):

its 3 times as big

OpenStudy (geekfromthefutur):

oh ok

OpenStudy (geekfromthefutur):

@thatk12kid000 3% 30% 300% 3000%

OpenStudy (geekfromthefutur):

@jim_thompson5910 C? or A?

jimthompson5910 (jim_thompson5910):

900/300 = 3 convert 3 to a percentage to get 3*1 = 3*100% = 300%

OpenStudy (geekfromthefutur):

Thank you that makes more sense

jimthompson5910 (jim_thompson5910):

yw